# Neutrinos as a new tool to characterise the Milky Way Centre

### v1

@@ Abstract
In conclusion, our work shows how neutrino observations will provide valuable insights into the gas distribution of the CMZ and the fraction of gamma-rays generated through leptonic and hadronic processes in the Galactic Centre. While achieving high-significance model discrimination may require exposure levels that will not be reached for another two decades, neutrino telescopes will ultimately provide a fully independent test of the gas density inferred from mass tracers, offering the most robust constraints in terms of systematic uncertainties. Improved measurements of gas properties will facilitate the development of star-formation models in the CMZ, and this knowledge can be generalised to modeling systems with similar environmental conditions, including galaxies beyond the Milky Way.
@@
Acknowledgements—The authors would like to thank Foteini Oikonomou, Kate Pattle, and Sheng-Jun Lin for useful discussions. This work has been supported by the Cosmoparticle Initiative of UCL. PCWL is supported by a UCL Graduate Research Scholarship and a UCL Overseas Research Scholarship. MA acknowledges support from the Science and Technology Facilities Council, part of the UK Research and Innovation (Grant No. ST/T004169/1). ERO is an international research fellow under the Postdoctoral Fellowship of the Japan Society for the Promotion of Science (JSPS), supported by JSPS KAKENHI Grant Number JP22F22327, and also acknowledges support from the RIKEN Special Postdoctoral Researcher Program for junior scientists.

### v2

@@ Abstract
While achieving high-significance model discrimination may require two decades or more depending on the assumptions, neutrino telescopes will ultimately provide a fully independent and solid test of the gas density inferred from mass tracers. This improved measurement of gas properties will facilitate the development of star-formation models in the CMZ, and this knowledge will be generalised to model galaxies beyond the Milky Way.
In addition, the next decade will witness a new wave of neutrino telescopes coming online alongside the Cherenkov Telescope Array Observatory (CTAO). CTAO will enable a step-change in the precision of gamma-ray measurements of the CMZ [33]. In particular, by detecting photons up to 300 TeV, CTAO will provide crucial insights into the presence of a possible cut-off affecting the neutrino spectra above the range explored by HESS and HAWC. At the same time, the next generation of neutrino telescopes will provide constraints on the gamma-ray flux due to hadronic processes, potentially leading to the identification of gamma-ray sources powered by leptonic processes. The reciprocal feedback between these two messengers will drive transformative progress and push the frontiers of multimessenger astronomy.
Acknowledgements—The authors would like to thank Foteini Oikonomou, Kate Pattle, and Sheng-Jun Lin for useful discussions.
We thank the referees for their valuable feedback.
We thank ESA for making Herschel data available, and NAOJ for making NRO data available.
We are also grateful to the PLEM group
(http://github.com/PLEnuM-group/Plenum)
for making their software open source.
@@
This work has been supported by the Cosmoparticle Initiative of UCL. PCWL is supported by a UCL Graduate Research Scholarship and a UCL Overseas Research Scholarship. MA acknowledges support from the Science and Technology Facilities Council, part of the UK Research and Innovation (Grant No. ST/T004169/1). ERO is an international research fellow under the Postdoctoral Fellowship of the Japan Society for the Promotion of Science (JSPS), supported by JSPS KAKENHI Grant Number JP22F22327, and also acknowledges support from the RIKEN Special Postdoctoral Researcher Program for junior scientists.